Hi,
Hello i am Vanessa my daughter is really having hard time with this and i would like to know what is the answer and how you got it so i could help her learn how to do it.My friend gave me this website just now and she and i dont know it please answer fast.
A number has 4 digits.
No digits in the number are repeated.
The digit in the tens place is three times the digit in the thousands place.
The number is odd.
The sum of the digits in the number is 27.
What is the number?

Answer by ankor@dixie-net.com(22740)

A number has 4 digits.
No digits in the number are repeated.
The digit in the tens place is three times the digit in the thousands place.
The number is odd.
The sum of the digits in the number is 27.
What is the number?
:
Let w = the thousands digit
Let x = the 100's
Let y = the 10's
Let z = the units
:
The digit in the tens place is three times the digit in the thousands place.
y = 3w
Use some an assumption and some logic here
With a total of 27, we have to assume one digit is 9, w has to be less than 4
let y = 9
then
w = 3
:
So we have:
3 + x + 9 + z = 27
12 + x + z = 27
x + z = 27 - 15
x + z = 15
there are only 2 single digit pairs that add up to 15
6 & 9
7 & 8
we have already used 9 (no repeats), it must be 7 & 8
the last digit has to be odd, therefore we have
x = 8
z = 7
:
Our number: 3897